OpenStudy (ayubie):

Find the linear approximation to f(x) = sqrt{25-x ^{2}} near 3.

myininaya (myininaya):

Linear approximation just means find the tangent line to f at x=3

OpenStudy (ayubie):

so all I need to do is find f'(3)?

myininaya (myininaya):

well that is the slope of the tangent line

myininaya (myininaya):

you need to find the tangent line

OpenStudy (ayubie):

oops, that's what I meant.

myininaya (myininaya):

y=f'(3)(x-3)+f(3)

OpenStudy (ayubie):

so for f'(3) I got 3/16, but what do I plug in for x in (x-3)?

OpenStudy (ayubie):

i mean 3/4, not 3/16.

myininaya (myininaya):

do you mean -3/4?

OpenStudy (ayubie):

oh my goodness. yes. finals are killing me, can you tell?

myininaya (myininaya):

so we have y=-3/4 *(x-3)+f(3)

myininaya (myininaya):

all you have to do now is find f(3)

myininaya (myininaya):

this is your linear approximation for values near x=3

myininaya (myininaya):

and I'm talking about y=-3/4 * (x-3)+f(3)

OpenStudy (ayubie):

okay, so x-3 will just stay in my answer?

myininaya (myininaya):

all i was doing was using the point slope formula to find the tangent line y-y_1=m(x-x_1) I just added y_1 on both sides y=m(x-x_1)+y_1 x_1 is 3 y_2 is f(3)

myininaya (myininaya):

if you want to distribute the -3/4 and combine like terms you may want to if you are entering answers into a machine